Find the argument of z = (6 + 5i)20 on the principal branch, i.e. Arg z.
Answer #1

Given 7= (6+52) 20 = 120+ 100€ =xtiy . =7 x = 120, y = 100 2. Arga= Tañi (H/x-) = Tani (120) = Taño ( 3 ) = 39.81°